OpenStudy (anonymous):

Why is i^i = e^(-pi/2)?

OpenStudy (anonymous):

i^(i*i)=i^(-1)=-i=cos(-pi/2)+isin(-pi/2)=e^(-ipi/2). taking ith root, i^i=e^(-pi/2)